# Multispectral airborne laser scanning for tree species classification: a benchmark of machine learning and deep learning algorithms

### v1

@@ Abstract
Climate-smart and biodiversity-preserving forestry demands precise information on forest resources, extending to the individual tree level. Multispectral airborne laser scanning (ALS) has shown promise in automated point cloud processing and tree segmentation, but challenges remain in identifying rare tree species and leveraging deep learning techniques. This study addresses these gaps by conducting a comprehensive benchmark of machine learning and deep learning methods for tree species classification. For the study, we collected high-density multispectral ALS data ($>1000$ $\mathrm{pts}/\mathrm{m}^{2}$) at three wavelengths using the FGI-developed HeliALS system, complemented by existing Optech Titan data (35 $\mathrm{pts}/\mathrm{m}^{2}$), to evaluate the species classification accuracy of various algorithms in a test site located in Southern Finland. We established a field reference dataset of over 6000 trees across nine species using a newly developed browser-based crowdsourcing tool, which facilitated efficient data annotation. The ALS data, including a training dataset of 1065 segments, was shared with the scientific community to foster collaborative research and diverse algorithmic contributions. Based on 5261 test segments, our findings demonstrate that point-based deep learning methods, particularly a point transformer model, outperformed traditional machine learning and image-based deep learning approaches on high-density multispectral point clouds. For the high-density ALS dataset, a point transformer model provided the best performance reaching an overall (macro-average) accuracy of 87.9% (74.5%) with a training set of 1065 segments and 92.0% (85.1%) with 5000 training segments. The best image-based deep learning method, DetailView, reached an overall (macro-average) accuracy of 84.3% (63.9%), whereas a random forest (RF) classifier achieved an overall (macro-average) accuracy of 83.2% (61.3%). For the sparser ALS dataset, an RF model topped the list with an overall (macro-average) accuracy of 79.9% (57.6%), closely followed by the point transformer at 79.6% (56.0%). Importantly, the overall classification accuracy of the point transformer model on the HeliALS data increased from 73.0% with no spectral information to 84.7% with single-channel reflectance, and to 87.9% with spectral information of all the three channels. By studying the classification accuracy as a function of point density, we conclude that the multispectral information seems especially important for sparse point clouds with 1–50 $\mathrm{pts}/\mathrm{m}^{2}$. Furthermore, we observed that the classification error follows a power law $\varepsilon(m)\approx m^{-\alpha}$ as a function of the training set size $m$ for both the point transformer model and a RF classifier, with the scaling exponent $\alpha$ being approximately twice higher for the point transformer. By extrapolating the fitted scaling laws, we estimate that a macro-average accuracy of 90% would require 14 000 training trees with the point transformer and $4.9\times 10^{6}$ training trees with the RF classifier, thus highlighting the advantage of deep learning approaches on large datasets.
@@
Keywords tree species, ALS, multispectral, deep learning, machine learning, benchmarking, point cloud
@@ 1 Introduction
Knowledge of the species-specific size distribution of forest stands is crucial for effective forest management planning and optimization of the wood supply chain. This information allows for decisions on tailored silvicultural treatments, such as harvesting schedules, thinning operations, and regeneration strategies. Reliable information on tree species also contributes to sustainable forest management, ensuring a continuous supply of high-quality timber products while maintaining ecological integrity. Namely, the diversity of tree species plays an important role in shaping forest ecosystems, and it affects factors such as productivity, resilience to disturbances, recovery rates, internal competition among trees, ecosystem health, economic potential and biodiversity. For example in the boreal forest zone, aspen (Populus tremula L.) is widely recognized as a keystone species, having an important role in supporting biodiversity (Kouki et al., 2004; Kivinen et al., 2020) since aspen stands often harbor a rich diversity of fungi, lichens, insects, birds, and mammals (Kuusinen, 1994; Angelstam and Mikusiński, 1994; Tikkanen et al., 2006; Remm et al., 2017).
@@
Although remote sensing techniques, such as satellite imagery and airborne laser scanning (ALS), have shown great promise in forest monitoring (Hyyppä et al., 2000; Næsset et al., 2004; Holmgren and Persson, 2004; Ørka et al., 2009), accurate species classification at the individual tree level remains a challenging task. For example, previous ALS studies in the boreal forest zone have demonstrated a high classification accuracy for the dominant pine and spruce trees (Holmgren and Persson, 2004; Ørka et al., 2012; Yu et al., 2017; Hakula et al., 2023), but the classification accuracy of deciduous trees has remained markedly lower especially for under-represented species (Ørka et al., 2012; Axelsson et al., 2018; Yu et al., 2017). In general, the classification accuracy degrades if the number of species is high (Heinzel and Koch, 2011; Prieur et al., 2021) or if there are closely related species, particularly those with similar spectral and structural characteristics. Therefore, further efforts are needed to improve species classification using remote sensing techniques since accurate species classification will be required to enable multifunctional forestry serving multiple ecosystem services optimally at the same time.
@@
Today, airborne laser scanning is the most prevalent technique for forest inventory in the boreal forest zone, with a strong emphasis on economically efficient forest resource utilization. The state of the art in species classification using ALS is summarized in several review papers (Chen et al., 2024; Michałowska and Rapiński, 2021; Wang et al., 2018; Koenig and Höfle, 2016; Fassnacht et al., 2016) and a recent benchmarking study using the FOR-species20K open dataset by Puliti et al. (2025). In general, the accuracy of species classification is affected by several factors, including the point cloud quality (e.g. point density, beam footprint, sensing direction), additional features available for classification (intensity at one or multiple wavelengths, echo types, waveform information), the chosen classification method (machine learning or deep learning), and the difficulty of the studied forest area (number and similarity of species, tree density, size distribution of trees). Until recently, ALS-based species classification has been implemented using hand-crafted features and traditional supervised machine learning (ML) classifiers, such as random forests (RF) (Ørka et al., 2012; Dalponte et al., 2012; Yu et al., 2017), support vector machines (SVM) (Ørka et al., 2012; Dalponte et al., 2012, 2014) or discriminant analysis (Holmgren and Persson, 2004; Ørka et al., 2009; Heinzel and Koch, 2011). Typical features for species classification have included point cloud metrics representing the geometry of a tree (e.g. crown shape, tree height, density and height distributions of points, trunk hits, echo ranks, point counts), backscatter and intensity statistics (e.g. intensity/amplitude, reflectance and waveform distributions), and temporal changes in the canopy (e.g. the difference of intensity or average canopy height between leaf-on and leaf-off seasons) (Holmgren and Persson, 2004; Ørka et al., 2009; Fassnacht et al., 2016; Koenig and Höfle, 2016; Lin and Hyyppä, 2016; Yu et al., 2017).
@@
During the past few years, deep learning (DL) approaches have been shown to outperform traditional machine learning methods in species classification (Xi et al., 2020; Liu et al., 2021). As a result, several deep learning methods have been recently proposed for species classification both using ALS data (Hell et al., 2022; Marinelli et al., 2022; Fan et al., 2023; Lin et al., 2024) and terrestrial laser scanning (TLS) data (Seidel et al., 2021; Liu et al., 2022a). Importantly, deep learning approaches can be divided into image-based 2D methods (Seidel et al., 2021; Marinelli et al., 2022) and point-based 3D methods (Xi et al., 2020; Liu et al., 2021; Fan et al., 2023; Lin et al., 2024). In image-based DL methods, the point cloud is projected into multiple views from different perspectives, which are given as an input for a 2D convolutional neural network (CNN), such as ResNet (He et al., 2016), DenseNet (Huang et al., 2017), or YOLO (Jocher et al., 2023). On the other hand, the input of point-based methods is the point cloud itself with potential additional features. Typical neural net architectures of point-based methods include, e.g., PointNet (Qi et al., 2017a), PointNet++ (Qi et al., 2017b), dynamic graph CNN (DGCNN) (Wang et al., 2019b), and point transformer (Zhao et al., 2021). Deep learning methods require large amounts of training data to reach their potential, which has resulted in a need for large open datasets enabling the development and benchmarking of new deep learning-based classification techniques. To this end, Puliti et al. (2025) recently presented FOR-species20K, a 33 species open dataset consisting of 20 000 trees collected using terrestrial, mobile and unmanned aerial vehicle based laser scanning with the goal to facilitate the development of sensor-agnostic classification approaches using only the point cloud geometry. Furthermore, Puliti et al. (2025) presented a comparison of seven sensor-agnostic deep learning methods concluding that the studied image-based deep learning models marginally outperformed the point-based methods.
@@
Backscatter intensity (or reflectance), even though hard to standardize across different sensors, has been successfully applied for species classification from the beginning of ALS research. Holmgren and Persson (2004) showed that a combination of geometric and intensity features led to an overall accuracy of 95% for the discrimination of spruce and pine. Since then intensity and geometric features have been used in many studies on tree species classification. Ørka et al. (2009) demonstrated a classification accuracy of 88% between Norway spruce and birch for about 350 trees, and Korpela et al. (2010) achieved an accuracy of 88–90% for Scots pine, Norway spruce, and birch across 13000 trees. Importantly, multiple studies have demonstrated that the classification accuracy obtained with geometric features can be significantly improved with the addition of intensity features (Axelsson et al., 2018; Yu et al., 2017; Lin and Hyyppä, 2016; Suratno et al., 2009; Shi et al., 2018a; Hakula et al., 2023; Ørka et al., 2009). For example, Ørka et al. (2009), Yu et al. (2017), and Hakula et al. (2023) reported 11, 9.9 and 13.5 percentage point improvements, respectively, to the overall classification accuracy when introducing single-channel intensity features to the classification. Motivated by the success of single-channel intensity features, the integration of ALS data with passive multispectral or even hyperspectral data has also been studied to further increase the classification accuracy (Holmgren et al., 2008; Ørka et al., 2012; Dalponte et al., 2014; Deng et al., 2016; Kamińska et al., 2018, 2021; Zhong et al., 2022; Quan et al., 2023; Lisiewicz et al., 2025).
Multispectral ALS using multiple laser wavelengths provides a further alternative to enhance the accuracy of species classification with potential to meet the increasing requirements for biodiversity reporting. Originally, multispectral laser scanning was developed to increase the capacity of ALS point clouds for automated object recognition (Kaasalainen et al., 2007; Hyyppä et al., 2013). More recently, multispectral ALS has been applied for tree species classification (Yu et al., 2017; Axelsson et al., 2018; Amiri et al., 2019; Prieur et al., 2021; Rana et al., 2022; Hakula et al., 2023; Wang et al., 2024), with some studies observing a reduction of the classification error by up to $\sim 30$% compared to the use of single-channel features (Axelsson et al., 2018; Amiri et al., 2019; Hakula et al., 2023). Despite the promise of deep learning for semantic segmentation on multispectral data (Reichler et al., 2024; Oinonen et al., 2024; Ruoppa et al., 2025), deep learning approaches for species classification have been studied to date in only a single paper focusing on sparse (30 pts/$\mathrm{m}^{2}$) multispectral ALS data (Wang et al., 2024). Furthermore, the paper by Hakula et al. (2023) is the only study utilizing high-density ($>1000$ pts/$\mathrm{m}^{2}$) multispectral ALS data for species classification. In Hakula et al. (2023), the use of multispectral ALS data reduced the overall classification error by 65% compared to the use of single-channel geometric features and by 31% compared to single-channel geometric and intensity features for the main tree species, including spruce, pine and birch. The use of deep learning methods for species classification on high-density multispectral ALS data has not been studied previously.
@@
Consequently, there is a pressing need for more in-depth research on multispectral airborne laser scanning for tree species classification, especially in the presence of a wide variety of species and with high-density point clouds. To facilitate this advancement, we undertook the following actions:
  • We collected high-density ALS data (1300 pts/$\mathrm{m}^{2}$) using a novel multispectral scanner system in an area located in Southern Finland where we have previously collected sparse multispectral ALS data (35 pts/$\mathrm{m}^{2}$) using the Optech Titan system (Fernandez-Diaz et al., 2016).
@@
  • We established a high-quality field reference dataset comprising approximately 6000 trees and nine species, including aspen, that is vital for biodiversity in the boreal forests. The field reference dataset was carefully acquired using a crowdsourcing tool facilitating the collection of large datasets required by deep learning models.
@@ 2 Materials and Methods
2.1 Test site
Figure 1: (a) The Espoonlahti test site is located in Southern Finland. (b) Orthophoto of the Espoonlahti test site together with collected tree segments colored by the species in the reference dataset. The coverage of the HeliALS data is shown with the light blue line. (c) A close-up of the orthophoto shown in (b).
The test site shown in Fig. 1 (center point approximately 60.1462°N 24.6587°E) is located in Espoonlahti in the city of Espoo, directly 20 km west of the center of Helsinki, on the southern coast of Finland. The test site was selected due to the large diversity of tree species growing in the area. A total of 20–30 different tree species can be found, including natural and planted trees both in built-up environment and in forests, as the area comprises different high-rise and low-rise residential areas, public buildings, a sports park and some recreational and non-recreational forest areas. The forests are natural forests and typically either pine-dominated rocky forests or mixed forests with coniferous and deciduous species. The mixed forests are relatively dense, while the rocky forests are sparser. Tree species classified in the study included pine, spruce, birch, maple, aspen, rowan, oak, linden and alder (see more information in Section 2.4). Pine, spruce, birch, aspen, rowan and alder are typically occurring natural tree species in the area, while most maples, oaks and lindens are planted trees in the built-up environment.
@@ 2.2 Measurement systems and acquisition of laser scanning data
Resulting point density 581 pts/$\mathrm{m}^{2}$ 175 pts/$\mathrm{m}^{2}$ 519 pts/$\mathrm{m}^{2}$
@@
Field of view $40^{\circ}$ $40^{\circ}$ $40^{\circ}$
@@
Resulting point density 11 pts/$\mathrm{m}^{2}$ 13 pts/$\mathrm{m}^{2}$ 11 pts/$\mathrm{m}^{2}$
@@
The HeliALS dataset was acquired on July 20 and 28, 2023, from an altitude of 100 m with leaf-on conditions. A helicopter, flying at 14 m/s, carried a system of three Riegl drone laser scanners operating in infrared, near-infrared and green wavelengths, namely a VUX-1HA ($\lambda=1550$ nm), a miniVUX-1DL ($\lambda=905$ nm), and a VQ-840-G ($\lambda=532$ nm). Positioning was achieved using a NovAtel (LITEF) ISA-100C inertial measurement unit (IMU), a NovAtel PwrPak7 Global Navigation Satellite System (GNSS) receiver, and a NovAtel (Vexxis) GNSS-850 antenna. The initial performance of the HeliALS system was presented in Hakula et al. (2023), which was the first study focusing on tree species classification from high-density multispectral ALS data. After the publication, the green channel data have been optimized by increasing the receiver aperture of the VQ-840-G scanner from 3 mrad to 6 mrad. The data were processed to point clouds at the FGI. Further specifications of the HeliALS system and data acquisition are provided in Table 1.
The Optech Titan system operates in three channels with infrared ($\lambda=1550$ nm), near-infrared ($\lambda=1064$ nm) and green wavelengths ($\lambda=532$ nm), respectively. The data were acquired on June 14, 2016, when trees were in full leaf (Karila et al., 2019). The acquisition was carried out in cooperation with TerraTec Oy (Helsinki, Finland). A fixed-wing aircraft and a flying altitude of 700 m were used. Data from the three channels were obtained as three separate point clouds. After basic processing by TerraTec Oy, the data were delivered to the FGI in the Finnish ETRS-TM35FIN coordinate system with height values in the Finnish N2000 height system. Further specifications of the Optech Titan system and data acquisition are presented in Table 2.

### v2

@@ Abstract
Climate-smart and biodiversity-preserving forestry demands precise information on forest resources, extending to the individual tree level. Multispectral airborne laser scanning (ALS) has shown promise in automated point cloud processing, but challenges remain in leveraging deep learning techniques and identifying rare tree species in class-imbalanced datasets. This study addresses these gaps by conducting a comprehensive benchmark of deep learning and traditional shallow machine learning methods for tree species classification. For the study, we collected high-density multispectral ALS data (>1000>1000 $\mathrm{pts}/\mathrm{m}^{2}$) at three wavelengths using the FGI-developed HeliALS system, complemented by existing Optech Titan data (35 $\mathrm{pts}/\mathrm{m}^{2}$), to evaluate the species classification accuracy of various algorithms in a peri-urban study area located in southern Finland. We established a field reference dataset of 6326 segments across nine species using a newly developed browser-based crowdsourcing tool, which facilitated efficient data annotation. The ALS data, including a training dataset of 1065 segments, was shared with the scientific community to foster collaborative research and diverse algorithmic contributions. Based on 5261 test segments, our findings demonstrate that point-based deep learning methods, particularly a point transformer model, outperformed traditional machine learning and image-based deep learning approaches on high-density multispectral point clouds. For the high-density ALS dataset, a point transformer model provided the best performance reaching an overall (macro-average) accuracy of 87.9% (74.5%) with a training set of 1065 segments and 92.0% (85.1%) with a larger training set of 5000 segments. With 1065 training segments, the best image-based deep learning method, DetailView, reached an overall (macro-average) accuracy of 84.3% (63.9%), whereas a shallow random forest (RF) classifier achieved an overall (macro-average) accuracy of 83.2% (61.3%). For the sparser ALS dataset, an RF model topped the list with an overall (macro-average) accuracy of 79.9% (57.6%), closely followed by the point transformer at 79.6% (56.0%). Importantly, the overall classification accuracy of the point transformer model on the HeliALS data increased from 73.0% with no spectral information to 84.7% with single-channel reflectance, and to 87.9% with spectral information of all the three channels. Furthermore, we studied the scaling of the classification accuracy as a function of point density and training set size using 5-fold cross-validation of our dataset. Based on our findings, multispectral information is especially beneficial for sparse point clouds with 1–50 $\mathrm{pts}/\mathrm{m}^{2}$. Furthermore, we observed that the classification error follows a power law $\varepsilon(m)\propto m^{-\alpha}$ as a function of the training set size $m$, and the classification error of the point transformer reduced significantly faster with increasing training set size compared to RF.
@@
Keywords Tree species, Airborne laser scanning, Multispectral, Deep learning, Machine learning, Lidar
Published in ISPRS Journal of Photogrammetry and Remote Sensing, 2026. Final version available at: https://doi.org/10.1016/j.isprsjprs.2026.01.031 . © 2026. This manuscript version is made available under the CC-BY-NC-ND 4.0 license https://creativecommons.org/licenses/by-nc-nd/4.0/
@@ 1 Introduction
Knowledge of the species-specific size distribution of forest stands is crucial for effective forest management planning and optimization of the wood supply chain. This information allows for decisions on tailored silvicultural treatments, such as harvesting schedules, thinning operations, and regeneration strategies. Reliable information on tree species also contributes to sustainable forest management, ensuring a continuous supply of high-quality timber products while maintaining ecological integrity. Namely, the diversity of tree species plays an important role in shaping forest ecosystems, and it affects factors such as productivity, resilience to disturbances, recovery rates, internal competition among trees, ecosystem health, economic potential and biodiversity. For example in the boreal forest zone, aspen (Populus tremula L.) is widely recognized as a keystone species, having an important role in supporting biodiversity (kouki2004long; kivinen2020keystone), since aspen stands often harbor a rich diversity of fungi, lichens, insects, birds, and mammals (kuusinen1994epiphytic; angelstam1994woodpecker; tikkanen2006red; remm2017multilevel). Additionally, accurate tree species identification in municipalities is required for achieving the targeted restoration and effective enhancement of biodiversity and canopy cover within urban ecosystems (EU_nature; GER_nature; FIN_nature).
@@
Although remote sensing techniques, such as satellite imagery and airborne laser scanning (ALS), have shown great promise in forest monitoring (hyyppa2000accuracy; naesset2004laser; holmgren2004identifying; orka2009classifying), accurate species classification at the individual tree level remains a challenging task. For example, previous ALS studies in the boreal forest zone have demonstrated a high classification accuracy for the dominant pine and spruce trees (holmgren2004identifying; orka2012simultaneously; yu2017single; hakula2023individual), but the classification accuracy of deciduous trees has remained markedly lower especially for under-represented species (orka2012simultaneously; axelsson2018exploring; yu2017single). In general, the classification accuracy degrades if the number of species is high (heinzel2011exploring; prieur2021comparison) or if there are closely related species, particularly those with similar spectral and structural characteristics. Therefore, further efforts are needed to improve species classification using remote sensing techniques since accurate species classification will be required to enable multifunctional forestry serving multiple ecosystem services optimally at the same time.
@@
Today, airborne laser scanning is the most prevalent technique for forest inventory in the boreal forest zone, with a strong emphasis on economically efficient forest resource utilization. The state of the art in species classification using ALS is summarized in several review papers (chen2024tree; michalowska2021review; wang2018review; koenig2016full; fassnacht2016review) and a recent benchmarking study using the FOR-species20K open dataset by puliti2025benchmarking. In general, the accuracy of species classification is affected by several factors, including the point cloud quality (e.g. point density, noise, geometric accuracy affected by beam footprint), additional features available for classification (intensity at one or multiple wavelengths, echo types, waveform information), the chosen classification method (machine learning or deep learning), and the difficulty of the studied forest area (number and similarity of species, tree density, size distribution of trees). Until recently, ALS-based species classification has been implemented using hand-crafted features and traditional shallow supervised machine learning (ML) classifiers, such as random forests (RF) (orka2012simultaneously; dalponte2012tree; yu2017single), support vector machines (SVM) (orka2012simultaneously; dalponte2012tree; dalponte2014tree) or discriminant analysis (holmgren2004identifying; orka2009classifying; heinzel2011exploring). Typical features for species classification have included point cloud metrics representing the geometry of a tree (e.g. crown shape, tree height, density and height distributions of points, trunk hits, echo ranks, point counts), backscatter and intensity statistics (e.g. intensity/amplitude, reflectance and waveform distributions), and temporal changes in the canopy (e.g. the difference of intensity or average canopy height between leaf-on and leaf-off seasons) (holmgren2004identifying; orka2009classifying; fassnacht2016review; koenig2016full; lin2016comprehensive; yu2017single). Throughout the manuscript, we refer to traditional shallow machine learning classifiers as machine learning approaches following the convention in several recent papers in tree species classification (xi2020see; mayra2021tree; hell2022classification). In contrast, we use the term deep learning method to refer to deep neural networks that learn their features.
@@
During the past few years, deep learning (DL) approaches have been shown to outperform traditional shallow machine learning methods in species classification (xi2020see; liu2021tree). As a result, several deep learning methods have been recently proposed for species classification both using ALS data (hell2022classification; marinelli2022approach; fan2023tree; lin2024pctrees) and terrestrial laser scanning (TLS) data (seidel2021predicting; liu2022tree). Deep learning approaches can be divided into image-based 2D methods (seidel2021predicting; marinelli2022approach) and point-based 3D methods (xi2020see; liu2021tree; fan2023tree; lin2024pctrees). In image-based DL methods, the point cloud is projected into multiple views from different perspectives, which are given as an input for a 2D convolutional neural network (CNN), such as ResNet (he2016deep), DenseNet (huang2017densely), or YOLO (Yolov82023). On the other hand, the input of point-based methods is the point cloud itself with potential additional features. Typical neural net architectures of point-based methods include, e.g., PointNet (qi2017pointnet), PointNet++ (qi2017pointnet++), dynamic graph CNN (DGCNN) (wang2019dynamic), and point transformer (zhao2021point). Deep learning methods require large amounts of training data to reach their potential, which has resulted in a need for large open datasets enabling the development and benchmarking of new deep-learning-based classification techniques. To this end, puliti2025benchmarking recently presented FOR-species20K, a 33 species open dataset consisting of 20 000 trees collected using terrestrial, mobile and unmanned aerial vehicle based laser scanning with the goal to facilitate the development of sensor-agnostic classification approaches using only the point cloud geometry. Furthermore, puliti2025benchmarking presented a comparison of seven sensor-agnostic deep learning methods concluding that the studied image-based deep learning models marginally outperformed the point-based methods.
@@
Backscatter intensity (or reflectance), even though hard to standardize across different sensors, has been successfully applied for species classification from the beginning of ALS research. holmgren2004identifying showed that a combination of geometric and intensity features led to an overall accuracy of 95% for the discrimination of spruce and pine. Since then intensity and geometric features have been used in many studies on tree species classification. orka2009classifying demonstrated a classification accuracy of 88% between Norway spruce and birch for about 350 trees, and korpela2010tree achieved an accuracy of 88–90% for Scots pine, Norway spruce, and birch across 13000 trees. Importantly, multiple studies have demonstrated that the classification accuracy obtained with geometric features can be significantly improved with the addition of intensity features (axelsson2018exploring; yu2017single; lin2016comprehensive; suratno2009tree; shi2018tree; hakula2023individual; orka2009classifying). For example, orka2009classifying, yu2017single, and hakula2023individual reported 11, 9.9 and 13.5 percentage point improvements, respectively, to the overall classification accuracy when introducing single-channel intensity features to the classification. Motivated by the success of single-channel intensity features, the integration of ALS data with passive multispectral or even hyperspectral data has also been studied to further increase the classification accuracy (holmgren2008species; orka2012simultaneously; dalponte2014tree; deng2016comparison; kaminska2018species; kaminska2021single; zhong2022identification; quan2023tree; lisiewicz2025comprehensive).
[…]
@@
Consequently, there is a need for more in-depth research on multispectral airborne laser scanning for tree species classification, especially in the presence of a wide variety of species and with high-density point clouds. Thus, the primary goal of our study is to carry out an international benchmark of tree species classification methods on multispectral ALS data. To facilitate this advancement, we undertook the following actions:
[…]
@@
  • We established a high-quality field reference dataset comprising 6326 segments and nine species, including aspen, that is vital for biodiversity in the boreal forests. The field reference dataset was carefully acquired using a crowdsourcing tool facilitating the collection of large datasets required by deep learning models.
@@ 2 Materials and Methods
2.1 Study area
Figure 1: (a) The Espoonlahti study area is located in southern Finland. (b) Orthophoto of the Espoonlahti study area together with collected tree segments colored by the species in the reference dataset. The coverage of the HeliALS data is shown with the light blue line. Areas A and B are separated by the orange line. (c) A close-up of the orthophoto shown in (b).
The study area (center point approximately 60.1462°N 24.6587°E) is located in Espoonlahti in the city of Espoo, directly 20 km west of the center of Helsinki, on the southern coast of Finland. The study area shown in Fig. 1 consists of two sub-regions, A and B, both of which were used for the current study. The study area can be characterized as a peri-urban site since it is partially covered by forests and partially by suburban neighborhoods. The area was selected due to the large diversity of tree species growing in the area compared to typical managed Finnish boreal forests with only a handful of species. On the study area, a total of 20–30 different tree species can be found, including natural and planted trees both in suburban environment and in forests, as the area comprises different residential areas, public buildings, a sports park and some recreational and non-recreational, unmanaged forest areas. The forests are natural forests and typically either pine-dominated rocky forests or mixed forests with coniferous and deciduous species. The mixed forests are relatively dense, while the rocky forests are sparser. Tree species classified in the study included pine (Pinus sylvestris), spruce (Picea sp.), birch (Betula sp.), maple (Acer platanoides), aspen (Populus tremula), rowan (Sorbus sp.), oak (Quercus robur), linden (Tilia sp.) and alder (Alnus sp.), see more information in Section 2.4. Pine, spruce, birch, aspen, rowan and alder are typically occurring natural tree species in the area, while most maples, oaks and lindens are planted trees in the environment.
@@ 2.2 Measurement systems and acquisition of laser scanning data
[…]
@@
Field of view $^{\circ} 4040^{\circ}$ $40^{\circ}$
@@
[…]
@@
The HeliALS dataset was acquired on July 20 and 28, 2023, from an altitude of 100 m with leaf-on conditions. A helicopter, flying at 14 m/s, carried a system of three Riegl drone laser scanners operating in infrared, near-infrared and green wavelengths, namely a VUX-1HA ($\lambda=1550$ nm), a miniVUX-1DL ($\lambda=905$ nm), and a VQ-840-G ($\lambda=532$ nm). Positioning was achieved using a NovAtel (LITEF) ISA-100C inertial measurement unit (IMU), a NovAtel PwrPak7 Global Navigation Satellite System (GNSS) receiver, and a NovAtel (Vexxis) GNSS-850 antenna. The initial performance of the HeliALS system was presented in hakula2023individual, which was the first study focusing on tree species classification from high-density multispectral ALS data. After the publication, the green channel data have been optimized by increasing the receiver aperture of the VQ-840-G scanner from 3 mrad to 6 mrad. The data were processed to point clouds at the FGI. Further specifications of the HeliALS system and data acquisition are provided in Table 1.
The Optech Titan system operates in three channels with infrared ($\lambda=1550$ nm), near-infrared ($\lambda=1064$ nm) and green wavelengths ($\lambda=532$ nm), respectively. The data were acquired on June 14, 2016, when trees were in full leaf (karila2019effect). The acquisition was carried out in cooperation with TerraTec Oy (Helsinki, Finland). A fixed-wing aircraft and a flying altitude of 700 m were used. Data from the three channels were obtained as three separate point clouds. After basic processing by TerraTec Oy, the data were delivered to the FGI in the Finnish ETRS-TM35FIN coordinate system with height values in the Finnish N2000 height system. Further specifications of the Optech Titan system and data acquisition are presented in Table 2.